OpenStudy (anonymous):

I want to build a right triangular garden on the side of my house. Find the sides of the triangle if the hypotenuse is 6 feet and the two sides are equal in length. please help :)

OpenStudy (anonymous):

Answer A=4.24 ft B=5.34 ft C=6.78 ft D=8.56 ft ??????

OpenStudy (rsadhvika):

hyp^2 = leg^2 + leg^2

OpenStudy (rsadhvika):

6^2 = 2leg^2 18 = leg^2 ? = leg

OpenStudy (rsadhvika):

use calculator

OpenStudy (anonymous):

okay thanx